can someone help me pls?
solve.
if m varies inversely as t and m=7 when t=3, find m when t=6

Answer by ewatrrr(24785)

 
Hi,
m varies inversely as t OR
m = k/t |m=7 when t=3
7 = k/3
21 = k
m = 21/t | t = 6 THEN m = 21/6 = 7/2